# Star Formation Histories of Dwarf Lenticular Galaxies

### v1

@@ 3.1 Luminosity and Mass Fraction
Figure 5 presents the mean luminosity and mass fractions of the stellar populations as a function of stellar age for dS0 galaxies. We applied sigma clipping to exclude galaxies deviating by more than from the mean. The number of rejected galaxies accounts for less than of the entire sample. It is convenient to use lookback time () to describe the SFHs of galaxies. One prominent characteristic of the SFHs of dS0 galaxies is the presence of multiple bursts of star formation. The first period of active star formation commenced with an explosive burst at log() and concluded at log(). There was a gradual decrease in star formation during this initial period. In contrast, the second period of active star formation exhibited a peak at log() and extended from log() to log(). Afterward, a brief quiescent phase followed. Subsequently, there was a resurgence of star formation activity at log() , albeit at significantly reduced intensity. This pattern, which includes a period of quenched star formation between log() and log() , represents a common feature observed in the SFHs of early-type dwarf galaxies within the LG (Tolstoy et al., 2009).
@@
It’s worth noting that approximately 60% of the stellar mass was produced during the first period of star formation, and more than 99.8% of the stellar mass was produced by the end of the second period. Thus, concerning stellar mass, star formation virtually ceased at log() . However, in terms of luminosity, star formation continues until recently to produce young massive stars. The SFHs of dS0 galaxies exhibit striking similarities to the SFHs of dEs (Seo & Ann, 2023). This suggests that there are commonalities in the star formation histories of these two classes of early-type dwarf galaxies.
While the general trend of star formation activity during the first period exhibits a gradual decrease, it is not strictly monotonic. There is a brief period of significantly reduced star formation immediately adter the initial explosive starburst, followed by a slight increase that leads to a peak at log() . This second peak within the first period is comparable to the distinct peak at log() in the second period.
@@
In terms of luminosity fractions, the initial burst of dS0 galaxies which contributes of the present luminosity is much weaker than the strongest burst at log() and the stars formed in the first period of star formation contribute of the present luminosity in total. However, we consider the first period as the primary phase of star formation because of the stellar mass was formed during this interval. Moreover, over 50% of the mass is contributed by stars older than approximately 10 Gyr. The oldest stars, which formed at log() , currently contribute roughly 25% of the total stellar mass and around 10% of the total luminosity. As described below, the oldest stars are predominantly metal-poor (Z=0.0004) stars.
@@
After the second period of star formation, the general star formation behavior is characterized by a period of quenched star formation until log() , followed by increased star formation activity leading up to the present time. However, the total mass contribution by stars younger than approximately 0.1 Gyr is less than 5%. The near absence of star formation between log() and log() has resulted in the lack of intermediate-age stars older than approximately 0.1 Gyr. This feature, marked by highly suppressed star formation in intermediate-age stars, is also observed in some early-type dwarfs in the LG (Weisz et al., 2014a), as well as in dSphs and dEs in the local universe (Seo & Ann, 2023). The suppression of star formation is believed to be caused by stellar feedback mechanisms such as supernova explosions and stellar winds. The cold gas within galaxies gets heated and is easily expelled into the halos of early-type dwarfs.
The resurgence of star formation around log() is believed to be primarily influenced by the accretion of cold intergalactic medium (IGM). It contributes of the present luminosity and less than 0.2% of the total stellar mass. This hypothesis gains support from the observed metallicity of stars formed immediately following the period of quenched star formation, which can be accounted for by assuming the accretion of metal-free IGM, as elaborated below.
@@ 3.2 Metallicity Dependent Mass Fraction
Figure 6 shows the mean mass fractions of the stellar populations of 148 dS0s, divided into 6 metallicity groups, as a function of stellar age. As described above for the derivation of mean luminosity and mass fractions, we applied a sigma clipping method to exclude outliers. The stellar metallicities we used are Z=0.0001, Z=0.0004, Z=0.004, Z=0.008, Z=0.02, and Z=0.05. The most pronounced feature of the SFHs of dS0s is the lack of extremely metal-poor (Z=0.0001) stars formed at log () . Among the stars formed at log () , which contribute of the present stellar mass, more than of them are metal-poor and the rest have metallicities of Z=0.004, 0.008, 0.02. and 0.05, each of which contributes at most. Stars formed from the initial burst at log () are also mostly () metal-poor stars. The absence of extremely metal-poor stars formed in the first and second periods of star formation is also observed in dSphs and dEs (Seo & Ann, 2023) which is assumed to be due to pre-enrichnent during the re-ionization era.
The chemical evolution of dS0 galaxies during the second period of star formation closely parallels that of dEs (Seo & Ann, 2023). Specifically, it commences with intermediate metallicities (Z=0.004 and Z=0.008) and concludes with the emergence of extremely metal-rich stars. Within the second period of star formation, the predominant metallicity is the intermediate metallicity of Z=0.008, accounting for more than half of the stars formed during this phase. This high prevalence of stars with Z=0.008 is attributed to the rapid enrichment process in dS0 galaxies.
Due to this swift enrichment, extremely metal-rich stars are produced in the latter stages of the second period of star formation. In fact, all generations of stars formed during the first and second periods of star formation exhibit some extremely metal-rich stars, although their contribution to the current stellar mass is relatively small, with the exception of stars formed at log() .
The occurrence of a quenched period of star formation in dS0 galaxies could be attributed to the explosive starbursts that took place during the initial phase of star formation, which expelled the remaining gas into the galactic halo. Over time, this gas, retained within the halo of dS0 galaxies, eventually falls back and leads to star formation after a period of dormancy. Similar characteristics have been observed in dEs as well (Seo & Ann, 2023).
@@ 4.1 Mean cSFHs
As have seen in Figures 5 and 6, the mean SFH of dS0 galaxies varies significantly since the beginning of star formation at Gyr ago. To understand the mass assembly histoy of dS0 galaxies more clearly we examine the cumulative star formation histories (cSFHS) of dS0s in detail below. Here again, we exaimne the mean cSFH of dS0 galaxies rather than cSFH of indivual galaxies because the cSFH of infividual galaxy is very diverse. We derived a mean cSFH of dS0 galaxies by applying 3 clippng. The mean cSFH could help in identifying overarching trends and patterns that apply broadly to this class of galaxies, shedding light on their evolutionary processes. For comparision with other early-type galaxies, dSphs and dEs which were studied by Seo & Ann (2023), we plotted the cSFHs of these types together in Figure 7.
The general behaviour of the cSFHs of dS0s is characterized by a rapid increase of cSFH due to bursts of star formation occurring at log() , contributing to of the present stellar mass. Subsequently, star formation activity significantly decreases resulting in slow increase until log() . There is no discernible increase in cSFH due to the almost complete quenching of star formation from log().
The cSFH of dEs is very similar to that of dS0s, while there is a significant difference in the cSFH of dSphs. Seo & Ann (2023) examined the SFHs of dSph and dE galaxies and conclude that the significant difference beween dSphs and dEs is related to the origin of these galaxies. The dSph galaxies are thought to be originated from the primordial galaxies while dE galaxies, at least significant fraction of them, are transformed from late-type galaxies. The similarity between dS0s and dEs suggests that dS0s are also likely to transformed from late-type galaxies. The morphological properties which are characterised by the presence of central lens component strongly supports the transformation origin of dS0 galaxies. It is worth to note that, VCC (Binggeli et al., 1985) and EVCC (Kim et al., 2014) do not distinguish dSphs from dEs. In contrast, as shown in Figure 7, the cSFHs of these two types are signicantly different. The similarity in cSFHs of dS0s and dEs suggests potential commonalities in their evolutionary processes and histories. This similarity and difference in the cSFHs of three sub-types of early-type dwarfs may provide insights into the formation and evolution of early-type dwarf galaxies within the broader context of galaxy evolution.
@@ 4.2.1 Morphology
There are two distinct morphological features associated with dS0 galaxies, outer spiral arms and nucleation. Outer spiral arms are unique features observed specifically in dS0 galaxies, while nucleation is a common trait among early-type dwarfs. Despite nucleation being a prevalent characteristic in early-type dwarfs (encompassing dS0s, dEs, and dSphs), the fraction of nucleated dS0 galaxies is notably smaller than that of dEs and dSphs, both of which exhibit nucleation in of their populations. This discrepancy is likely due to the presence of the central lens component, which distinguishes dS0 galaxies as a distinct sub-type within the category of early-type dwarfs. The fraction of dS0 galaxies that possess outer spiral arms is 0.2.
Figure 8 illustrates the mean cSFHs of dS0 galaxies as a function of lookback time, categorized based on the presence or absence of outer spiral arms and nucleation. In CVCG, galaxies with outer spiral arms are classified as dS0, while those with nucleation are labeled as dS0. In the left panel of Figure 8, we present the mean cSFHs of these 30 dS0 galaxies along with the 118 remaining dS0 galaxies obtained via 3 clipping. The analysis reveals no significant difference in the cSFHs between dS0 galaxies and dS0 galaxies, with only a marginal difference between the two samples noted between log() = 10 and log() . Given the lack of significant differences between the cSFHs of dS0 galaxies and dS0 galaxies, we combine the two samples and refer to them collectively as dS0s for subsequent analysis.
Another important morphological characteristic that distinguishes early-type dwarf galaxies is the presence of nucleation. In the case of dwarf lenticular galaxies listed in Ann et al. (2015), of them exhibit nucleation, with significant variations in size and brightness observed among these galaxies. In the right panel of Figure 8, we present the mean cSFHs of dS0 galaxies, divided into those with and without nucleation. Similar to the cSFHs of dS0 galaxies categorized by the presence or absence of spiral arm features, the cSFHs of dS0 galaxies with and without nucleation are quite similar. However, a slightly more rapid star formation phase is observed between log() and log() in the cSFH of dS0 galaxies that lack nucleation.
The analysis reveals only a slight difference in cSFHs between log() and log() . Consequently, the presence or absence of additional morphological features, such as outer spiral arms and nucleation does not significantly influence the cSFHs of dS0 galaxies.
@@ 4.2.2 Stellar Mass
where and are the effective radius and the fiber radius, respectively, and f(r) is the Sérsic profile (Sérsic, 1968). We assumed an axis-symmetric luminosity distribution and used the fact that the luminosity within is half of the total luminosity. We use the and Sérsic index () that were determined by Seo & Ann (2022). For galaxies with unknown and , we use the mean and derived for dS0s. The number of galaxies with unknown Sérsic parameters is of the sample galaxies. The stellar masses of dS0 galaxies were found to fall within the range of M to M.
@@
Figure 9 shows a comparison between the present estimates of the total stellar mass and those from SDSS DR12. The total stellar masses provided by SDSS DR12 were derived from multi-band photometric images following Kauffmann et al. (2003) and Brinchmann et al. (2004). Approximately of the sample galaxies are missing in the stellar mass table of SDSS DR12. There is a strong correlation between the two sets of stellar masses. The tight correlation for the majority of sample galaxies is remarkable because the two approaches are completely independent.
@@
Given the belief that the mass of a galaxy plays a crucial role in star formation, particularly in early-type dwarf galaxies, we divided the dS0 galaxies into two groups based on their stellar mass. The high mass group comprises the dS0 galaxies with stellar masses greater than the median mass plus one standard deviation (), while the low mass group includes the dS0 galaxies with stellar masses smaller than the median mass minus one standard deviation ().
@@
In Figure 10, we present the mean cSFHs of dS0 galaxies along with the boundaries, segregated into these two groups, as a function of lookback time. We calculated the mean cSFHs by applying 3 clipping. As depicted in Figure 10, the cSFHs of dS0 galaxies exhibit significant differences when grouped by stellar mass. The average cSFH of the high mass group shows a more rapid increase compared to that of the low mass group. Notably, the cSFH of the low mass group lies outside the boundary of the high mass group for lookback times between log() and log() . This suggests a strong stellar mass dependence in the cSFHs of dS0 galaxies, which appears to be slightly more pronounced than the dependencies observed in dSphs and dEs, as reported in Seo & Ann (2023).
4.2.3 Colors
The selection of u-r colors as a representative color parameter for galaxies is based on the understanding that, among the various colors defined in the SDSS photometric system, u-r colors are most closely linked to the SFHs of galaxies. In Figure 11, we present the cSFHs of dS0 galaxies, categorized by their u-r colors. Similar to our approach for stellar mass, we divided galaxies into two color groups using the median and standard deviation () of u-r colors. Galaxies with u-r colors falling below the median minus were placed in the "blue" color group, while those with u-r colors exceeding the median plus were classified as "red" galaxies.
As depicted in Figure 11, there is a substantial difference in the cSFHs of the two color groups. Once again, the mean cSFH of blue dS0 galaxies lies outside the 1 boundary of the cSFHs of red dS0 galaxies. After the earliest period of star formation, red galaxies exhibit a more rapid formation of stars compared to blue galaxies. Red galaxies conclude their star formation before a lookback time of log() = 9.2, while blue galaxies continue forming stars until log() = 8. This suggests that the red colors observed in some dS0 galaxies are a result of the early quenching of star formation. The strong color dependence of cSFHs in dS0 galaxies appears to be a natural consequence, as star formation significantly influences the colors of galaxies.

### v2

@@ 3.1 Luminosity and Mass Fraction
Figure 5 shows the average luminosity and mass fractions of stellar populations in dS0 galaxies as a function of stellar age. We applied sigma clipping to the luminosity and mass fractions, excluding galaxies that deviated more than 3 standard deviations from the mean. Fewer than 3% of the galaxies were excluded. The resolution of stellar ages in Figure 5 reflects the age grid used in STARLIGHT. The age grid intervals are not uniform, ranging mostly from 0.1 dex and 0.2 dex with smaller interval for old isochrones. As previously mentioned, the age resolution is at best 0.1 dex and can be worse for spectra with less than 30.
@@
One prominent characteristic of the SFHs of dS0 galaxies is the presence of multiple bursts of star formation. The first period of active star formation began with an explosive burst at Gyr ago and concluded around 6 Gyr ago. There was a gradual decrease in star formation during this first period from 14 Gyr ago to 5 Gyr ago. In contrast, the second period of active star formation exhibited a peak at 2.5 Gyr ago, spanning from around 6 Gyr ago to around 1 Gyr ago. Subsequently, star formation activity significantly decreased, reaching a nearly quiescent state. This pattern, characterized by initial bursts followed by a second peak and then near-quiescent star formation, is prevalent among dS0 galaxies, with 70% exhibiting the initial burst and 85% the second peak. Intriguingly, a minor resurgence of star formation activity occurred at 0.1 Gyr ago, albeit at significantly reduced intensity. Also, the period of quenched star formation between 1 Gyr ago and 0.1 Gyr ago, represents a common feature observed in the SFHs of early-type dwarf galaxies within the LG (Tolstoy et al., 2009).
@@
Notably, around 60% of the stellar mass formed during the first star formation period, and by the end of the second period, this value exceeded 99.8%. Consequently, star formation in terms of stellar mass essentially terminated at roughly 1 Gyr ago. However, star formation continues to produce young, massive stars at a low level, as evidenced by the luminosity. The SFHs of dS0 galaxies exhibit striking similarities to the SFHs of dEs (Seo & Ann, 2023). This suggests that there are commonalities in the star formation histories of these two classes of early-type dwarf galaxies.
@@
In terms of luminosity fractions, the initial burst of dS0 galaxies which contributes 10% of the present luminosity is much weaker than the strongest burst at 2.5 Gyr ago and the stars formed in the first period of star formation contribute 30% of the present luminosity in total. However, we consider the first period as the dominant phase of star formation because it produced around 60% of the stellar mass. Furthermore, over half of the total stellar mass originates from stars older than roughly 10 Gyr. The oldest stars, formed around 14 Gyr ago, currently comprise about 27% of the total stellar mass and contribute around 10% of the total luminosity. As described below, these oldest stars are primarily metal-poor (Z=0.0004).
@@
Following the second star formation period, dS0 galaxies enter a phase of quenched star formation that lasts until roughly 0.1 Gyr ago. This is followed by a renewed increase in star formation activity that continues to the present day. However, the stars formed during this recent period contribute less than 0.3% to the total stellar mass, with a total luminosity fraction of about 20%. The near-complete absence of star formation between roughly 1 Gyr ago and 0.1 Gyr ago explains the lack of intermediate-age stars older than 0.1 Gyr. This suppression of star formation, characterized by a dearth of intermediate-aged stars, is also observed in some early-type dwarfs in the LG (Weisz et al., 2014a), as well as in dSphs and dEs in the local universe (Seo & Ann, 2023). Stellar feedback mechanisms, such as supernova explosions and stellar winds, are believed to be responsible for quenching star formation. These processes heat the cold gas within the galaxies, making it easier to be expelled into the halos of early-type dwarfs.
@@ 3.2 Metallicity Dependent Mass Fraction
Figure 6 presents the average mass fractions of stellar populations within 148 dS0 galaxies, categorized into six metallicity groups, as a function of stellar age. Similar to Figure 5, we excluded outliers using sigma clipping before calculating average mass fractions. The metallicities considered are Z=0.0001, Z=0.0004, Z=0.004, Z=0.008, Z=0.02, and Z=0.05. The most pronounced feature of the dS0 SFHs is the dearth of extremely metal-poor stars (Z=0.0001) formed before 10 Gyr ago. Among the stars formed before this time, contributing 50% of the current stellar mass, over 80% are metal-poor. The remaining stars have metallicities of Z=0.004, 0.008, 0.02. and 0.05, each contributibg at most 5%. Interestingly, stars formed during initial burst at 14 Gyr ago are also predominantly (80%) metal-poor stars. This lack of extremely metal-poor stars formed during the first star formation period is also observed in dSphs and dEs (Seo & Ann, 2023). This is thought to be a consequence of pre-enrichment during the re-ionization era.
The chemical evolution in dS0 galaxies during the second period of star formation closely parallels that of dEs (Seo & Ann, 2023). Specifically, it starts with the formation of intermediate metallicities (Z=0.004 and Z=0.008) and ends with the emergence of extremely metal-rich stars. Within the second period of star formation, the predominant metallicity is the intermediate metallicity of Z=0.008, accounting for more than half of the stars formed during this phase. This high prevalence of stars with Z=0.008 is attributed to the rapid enrichment process in dS0 galaxies.
Due to this swift enrichment, extremely metal-rich stars are produced in the latter stages of the second period of star formation. In fact, all generations of stars formed during the first and second periods of star formation exhibit some extremely metal-rich stars, although their contribution to the current stellar mass is relatively small, with the exception of stars formed at 1.2 Gyr ago.
The occurrence of reduced star formation in dS0 galaxies after initial burst of star formation could be attributed to the explosive starbursts that took place during the initial phase. The tremendous energy released during the starbursts efficiently expelled the remaining gas into the galactic halo. However, this gas is not lost forever. Over time, some of the gas retained within the halo, eventually falls back, leading to star formation after 6 Gyr ago, i.e., initiating a second period of star formation. Similar behavior has also been observed in dEs as well (Seo & Ann, 2023).
@@ 4.1 Mean cSFHs
As shown in Figures 5 and 6, the mean SFH of dS0 galaxies varies significantly since the beginning of star formation at the lookback time of 14 Gyr. To understand the mass assembly history of dS0 galaxies more clearly we examine the cumulative star formation histories (cSFHs) of dS0s in detail below. We examine the mean cSFH of dS0 galaxies rather than cSFH of individual galaxies because, as shown in the appendix, the cSFH of individual galaxy is very diverse. We derived a mean cSFH of dS0 galaxies by applying 3 clipping. The mean cSFH could help to identify overarching trends and patterns that apply broadly to this class of galaxies, shedding light on their evolutionary processes. For comparison with other early-type galaxies, dSphs and dEs which were studied by Seo & Ann (2023), we plotted the cSFHs of these types together in Figure 7.
The general behaviour of the cSFHs of dS0s is characterized by a rapid increase of cSFH due to bursts of star formation occurring before the lookback time of 10 Gyr, contributing to 50% of the present stellar mass. Subsequently, star formation activity significantly decreases resulting in slow increase until the lookback time of 0.3 Gyr. There is no discernible increase in cSFH due to the almost complete quenching of star formation from the lookback time of 0.3 Gyr.
The cSFHs of dEs closely resemble those of dS0s, while dSphs exhibit significantly different cSFH patterns. Seo & Ann (2023) attributed the difference in the cSFHs of dEs and dSphs to their origins. The majority of dSphs are primordial galaxies while dEs, at least significant fraction of them, are transformed from late-type galaxies. The similarity between dS0s and dEs suggests that dS0s may also originate from late-type galaxies. The morphological properties which are characterised by the presence of central lens component strongly supports the transformation origin of dS0 galaxies. It is worth to note that, VCC (Binggeli et al., 1985) and EVCC (Kim et al., 2014) do not differentiate between dSphs and dEs. However, as shown in Figure 7, the cSFHs of these two types are signicantly different. The similarity in cSFHs between dS0s and dEs suggests potential commonalities in their evolutionary history. This contrast and comparison between the cSFHs of three sub-types of early-type dwarfs may provide insights into the formation and evolution of early-type dwarf galaxies within the broader context of galaxy evolution.
@@ 4.2.1 Morphology
There are two distinct morphological features associated with dS0 galaxies, outer spiral arms and nucleation. Outer spiral arms are unique features observed specifically in dS0 galaxies, while nucleation is a common trait among early-type dwarfs. Despite nucleation being a prevalent characteristic in early-type dwarfs (encompassing dS0s, dEs, and dSphs), the fraction of nucleated dS0 galaxies is smaller than those of dEs and dSphs, both of which exhibit nucleation in 85% of their populations. This discrepancy is likely due to the presence of the central lens component, which distinguishes dS0 galaxies as a distinct sub-type within the category of early-type dwarfs. The fraction of dS0 galaxies that possess outer spiral arms is 0.2.
Figure 8 illustrates the mean cSFHs of dS0 galaxies as a function of lookback time, categorized based on the presence or absence of outer spiral arms and nucleation. In the CVCG, dS0 galaxies with outer spiral arms are classified as dS0, while those with nucleation are labeled as dS0. In the left panel of Figure 8, we present the mean cSFHs of these 30 dS0 galaxies along with the 118 remaining dS0 galaxies obtained via 3 clipping. There is no significant difference in the cSFHs between dS0 galaxies and dS0 galaxies. A marginal difference exists between the two samples only between the lookback time of 10 Gyr and 2.5 Gyr. Given the lack of significant differences between the cSFHs of dS0p galaxies and dS0 galaxies, we will combine the two samples and refer to them collectively as dS0s in the subsequent analysis. In the right panel of Figure 8, we present the mean cSFHs of dS0s, divided into those with and without nucleation. Similar to the cSFHs of dS0 galaxies categorized by the presence or absence of spiral arm features, the cSFHs of dS0 galaxies with and without nucleation are quite similar. However, a slightly more rapid star formation phase is observed between the lookback time of 10 Gyr and 1 Gyr in the cSFH of dS0 galaxies that lack nucleation. Consequently, the presence or absence of additional morphological features, such as outer spiral arms and nucleation does not significantly influence the cSFHs of dS0 galaxies.
@@ 4.2.2 Stellar Mass
where and are the effective radius and the fiber radius, respectively, and f(r) is the Sérsic profile (Sérsic, 1968). We assumed an axis-symmetric luminosity distribution and used the fact that the luminosity within is half of the total luminosity. We use the and Sérsic index () that were determined by Seo & Ann (2022). For galaxies with unknown and , we use the mean and derived for dS0s. The average and it ranges from 0.05 to 0.6. The number of galaxies with unknown Sérsic parameters is of the sample galaxies. The stellar masses of dS0 galaxies were found to fall within the range of 5 M⊙ to M⊙.
@@
Figure 9 shows a comparison between the present estimates of the total stellar mass and those from SDSS DR12. The total stellar masses provided by SDSS DR12 were derived from multi-band photometric images following Kauffmann et al. (2003) and Brinchmann et al. (2004). Approximately of the sample galaxies are missing in the stellar mass table of SDSS DR12. There is a pretty good correlation between the two sets of stellar masses because they appear to scatter around the one-to-one relation. The rms in the correlation is dex, which is larger than that of the correlation in the age estimates, dex. The rms error of 0.4 dex is in a good agreement with the rms reported by Cid Fernandes et al. (2005). In the derivation of rms error we exclude outliers, indicated by ’x’ symbols (cyan) in Figures 4 and 9.
@@
Given the current consensus in the community that the mass of a galaxy plays a crucial role in star formation, particularly in early-type dwarf galaxies, we divided the dS0s into two groups based on their stellar mass. The high mass group comprises the dS0s with stellar masses greater than the median mass plus one standard deviation (), while the low mass group includes the dS0s with stellar masses smaller than the median mass minus one standard deviation ().
@@
In Figure 10, we present the mean cSFHs of dS0s along with the boundaries, segregated into these two groups, as a function of lookback time. We calculated the mean cSFHs by applying 3 clipping. As depicted in Figure 10, the cSFHs of dS0s exhibit significant differences when grouped by stellar mass. The average cSFH of the high mass group shows a more rapid increase compared to that of the low mass group. Notably, the cSFH of the low mass group lies outside the boundary of the high mass group for lookback times between 3 Gyr and 1.6 Gyr. This suggests a strong stellar mass dependence in the cSFHs of dS0s, which appears to be slightly more pronounced than the dependencies observed in dSphs and dEs, as reported in Seo & Ann (2023).